Shift report 20010819 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Aug 19 2001 - 07:56:38 EDT

  • Next message: Apache: "Shift report 20010819 08:00-16:00"

    ------------------------------------------------------------
    Date: Sun Aug 19 07:56:38 2001, Shift: 00:00-08:00
    Supervisor: sjs, with: am
    Log book#7 pages 128 - 131
    
    Shift Summary:
    ------------------------------------------------------------
    No beam until 1:18
    Run 4284 started at 1:45 after raising voltages. A polarity.
    At 3:45:  "Blue quench link interlock".  End run 4284.
    106K trigger 1 events.
    
    No running during last half of shift due to 
    machine problems:..."AtR permit problem"...
    "Permit Link problem"...
    "Performing hysteresis ramp".....end of shift.
    
    Total running time during shift: about 2 hours.
    
    Scan of supermon spectra indicate the following potential
    problems:  
    >About a dozen consecutive RHIC ADC channels missing (bad
    ADC or paddle card suggested by RD)
    >H1 TDC bot 33 missing
    >TOFW ADC top 49 missing
    >TOFW TDC Bot 50 missing
    >About a dozen, scattered signals missing in T3, T4, and T5
    X,Y,U,V spectra.
    
    Some of these issues have been previously noted.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Aug 19 2001 - 07:57:33 EDT